• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name spiegel

    Log entry time 22:42:40 on July24,2006

    Entry number 175158


    Determined lead flows A & B had droped to 60~61 SLM on Left Q2. Increased flows to 68 on both leads. Reset main contactor (manual lever) cleared interlock faults. Had problem switching polarity of magnet in remote mode. Changed to negitive polaritry in manual mode and returned magnet to remote control. Ramped to 100A in remote.